OpenStudy (anonymous):

ESTIMATE the difference between 11/12 and 4/9. Also, while you answer can you write or explain me the way you did it?. Thank you.

OpenStudy (anonymous):

\[\frac{ 11 }{ 12 } \frac{ 4 }{ 9 } ESTIMATED \]

OpenStudy (anonymous):

You could say 11/12 is close to 1 and 4/9 is close to 1/2, so the difference is around 1-1/2 = 1/2.

OpenStudy (anonymous):

Thank you very much.